Solutions of the inequality [(4x2-3x-1)^1/2]>= 2x-3 are?


1
Expert's answer
2021-09-07T00:53:44-0400

"[(4x^2-3x-1)^{\frac{1}{2}}] ≥ 2x-3"


"\sqrt{(4x^2-3x-1)} ≥ 2x-3"


Determine the defined range


 "\sqrt{4x^2-3x-1} ≥ 2x-3

,x\in[ ∞,-\frac{1}{4}]
 ∪
[1,+∞]"


Separate the inequality into 2 possible cases


"\sqrt{(4x^2-3x-1)} ≥ 2x-3 ," "2x-3 \n \u2265 0"


"\sqrt{(4x^2-3x-1)} ≥ 2x-3 ," "2x-3 \n <0"


Solve the inequality for x


"x≥\frac{10}{9},2x-3
≥0"


"\sqrt{(4x^2-3x-1)} ≥ 2x-3 , 2x-3<0"


"x≥\frac{10}{9},
x≥\frac{3}{2}"


"\sqrt{(4x^2-3x-1)} ≥ 2x-3 , 2x-3<0"


Since the left hand side is always positive or zero, and the right hand side is always negative, the statement is true for any value of x


"x≥\frac{10}{9},
x≥\frac{3}{2}"


"x\in \reals, 2x-3<0"


"x\in \reals, x<\frac{3}{2}"


Find the intersection

"x\in [\frac{3}{2},+∞]"


"x\in \reals, x<\frac{3}{2}"


"x\in [-∞,\frac{3}{2}]"


Find the union

"x\in \reals
, x\in[ -∞,-\frac{1}{4}]
 ∪

[1,+∞]"


Find the intersection of the defined solution and the defined range


"x\in[ -∞,-\frac{1}{4}]
∪

[1,+∞]"